# Stability and Decay for the 2D Anisotropic Navier-Stokes Equations with Fractional Horizontal Dissipation on $\mathbb{R}^{2}$

### v1

@@ Abstract
The stability problem for the 2D Navier-Stokes equations with dissipation in only one direction on $\mathbb{R}^{2}$ is not fully understood. This dissipation is in the intermediate regime between the fully dissipative Navier-Stokes and the inviscid Euler. Navier-Stokes solutions in $\mathbb{R}^{2}$ decay algebraically in time while Euler solutions can grow rather rapidly in time. This paper solves the fundamental stability and large-time behavior problem on the anisotropic Navier-Stokes with fractional dissipation $\Lambda_{1}^{2s}$ for all $0\leq s<1$. The case $s=1$ corresponds to the standard one directional dissipation $\partial_{1}^{2}$. Different techniques are developed to treat different ranges of fractional exponents: $0\leq s\leq\frac{3}{4}$, $\frac{3}{4}<s<\frac{11}{12}$, and $\frac{11}{12}\leq s<1$. The final range is the most difficult case, for which we introduce the spatial polynomial $A_{2}$ weights and exploit the boundedness of Riesz transforms on weighted $L^{2}$-spaces.
@@ 1 Introduction
$u_{0}=(u_{0,1},u_{0,2})=\nabla^{\perp}\psi_{0}:=(-\partial_{2}\psi_{0},\partial_{1}\psi_{0}),$
@@
$\|u_{0}\|_{H^{k}}+\|\Lambda_{1}^{-\sigma}\psi_{0}\|_{L^{2}}+\|\Lambda_{1}^{-\sigma}u_{0,1}\|_{L^{2}}+\|\Lambda_{1}^{-\sigma}\partial_{2}u_{0,1}\|_{L^{2}}\leq\varepsilon.$ (1.4)
@@
$\|\Lambda_{1}^{-\sigma}u(t)\|_{H^{k}}+\nu^{\frac{1}{2}}\left(\int_{0}^{t}\|\Lambda_{1}^{s-\sigma}u(\tau)\|_{H^{k}}^{2}\,d\tau\right)^{\frac{1}{2}}\leq C\varepsilon,\qquad\forall\,t\geq 0,$
@@
A special feature of this result is that the second velocity component $u_{2}$ decays faster than the first component $u_{1}$. This anisotropic decay is revealed through the use of the stream function and is a key ingredient in the proof of the theorem. The enhanced decay of $u_{2}$ plays a crucial role in closing the nonlinear estimates. Moreover, since the Navier-Stokes equations do not propagate negative Sobolev regularity of the form $\Lambda^{-\sigma}u$ for $\sigma\geq\tfrac{1}{2}$, the restriction $\sigma<\tfrac{1}{2}$ is necessary.
@@
The derivation of the above estimates parallels that of Theorem 1.2, we only present the estimates for $u_{2}$, which differs slightly, in Appendix B.
Once again, these estimates shows that the second velocity component $u_{2}$ exhibits a faster decay rate than $u_{1}$, reflecting the underlying anisotropic structure of the system.
@@
$\|\Lambda_{1}^{-\sigma}u(t)\|_{H^{k}}+\nu^{1/2}\left(\int_{0}^{t}\|\Lambda_{1}^{s-\sigma}u(\tau)\|_{H^{k}}^{2}\,d\tau\right)^{1/2}<\varepsilon,\qquad\forall\,t\geq 0.$
@@ 2 Anisotropic Estimates and Tools
A weight function $d\colon\mathbb{R}^{d}\to[0,\infty)$ is said to belong to the $A_{p}$ class if:
@@
    $d\in L_{\mathrm{loc}}^{1}(\mathbb{R}^{d})$ and $d(x)>0$ almost everywhere;
@@
    $\left(\frac{1}{|Q|}\int_{Q}d(y)\,dy\right)\left(\frac{1}{|Q|}\int_{Q}d(y)^{-\frac{1}{p-1}}\,dy\right)^{p-1}\leq C,\qquad 1<p<\infty,$
@@
$N_{1}:=-2\sum_{i,j=1}^{2}\sum_{\alpha=1}^{k}\binom{k}{\alpha}\int_{\mathbb{R}^{2}}\partial_{1}^{\alpha}u_{j}\,\partial_{j}\partial_{1}^{k-\alpha}u_{i}\,\partial_{1}^{k}u_{i}\,dx,$
@@
$N_{2}:=-2\sum_{i=1}^{2}\sum_{\alpha=1}^{k}\binom{k}{\alpha}\int_{\mathbb{R}^{2}}\partial_{2}^{\alpha}u_{1}\,\partial_{1}\partial_{2}^{k-\alpha}u_{i}\,\partial_{2}^{k}u_{i}\,dx,$
@@
$=$  $\;2\sum_{i=1}^{2}\sum_{\alpha=1}^{k}\binom{k}{\alpha}\int_{\mathbb{R}^{2}}\partial_{1}\partial_{2}^{\alpha-1}u_{1}\,\partial_{2}\partial_{2}^{k-\alpha}u_{i}\,\partial_{2}^{k}u_{i}\,dx,$
@@
For each $\ell\in\{1,2\}$, we decompose the analysis of $F^{III}_{H}$ into two parts.
@@
$+\sum_{\beta=1}^{k-1}\int_{\mathbb{R}^{2}}\varphi(\xi_{1})\mathcal{F}_{x_{1}}\big(\partial^{k}_{\ell}f\partial^{k}_{\ell}f\big)(\xi_{1},x_{2})\mathcal{F}_{x_{1}}\big(\partial_{1}f\big)(\xi_{1},x_{2})d\xi_{1}dx_{2}$
@@
In this section, we prove Theorem 1.2. The argument is based on a bootstrap approach combined with refined energy estimates that yield global-in-time decay and stability.
@@
$\|\Lambda_{1}^{-\sigma}u(t)\|_{H^{k}}$ $\leq\tfrac{1}{2}C_{0}\varepsilon,$ (4.7)
@@
4.1 Estimate for $\|\Lambda_{1}^{-\sigma}u\|_{L^{2}}$
$:=$ $M_{0}.$
Integrating in time on equation (4.13), we obtain
By Lemma 2.6,
Using the divergence-free condition of $u$, we have
It can be verified that
$s<\tfrac{5}{12}+\sigma<\tfrac{-18\sigma^{2}+25\sigma+3}{12}.$
As a consequence, we obtain
$\int_{0}^{t}M_{0}(\tau)\,d\tau\leq CC_{0}^{3}\varepsilon^{3}.$
4.2 Estimate for $\sum_{m=1}^{2}\|\Lambda_{1}^{-\sigma}\partial_{m}^{k}u\|_{L^{2}}$
Applying $\Lambda_{1}^{-\sigma}\partial_{m}^{k}$ to the first equation in (1.1), and then taking the $L^{2}$ inner product with $\Lambda_{1}^{-\sigma}\partial_{m}^{k}u$, we have
$\frac{d}{dt}\sum_{m=1}^{2}\left\|\Lambda_{1}^{-\sigma}\partial_{m}^{k}u\right\|_{L^{2}}^{2}+2\nu\sum_{m=1}^{2}\left\|\Lambda_{1}^{-\sigma}\Lambda_{1}^{s}\partial_{m}^{k}u\right\|_{L^{2}}^{2}$
$=$ $-2\sum_{m=1}^{2}\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}\partial_{m}^{k}(u\cdot\nabla u)\cdot\Lambda_{1}^{-\sigma}\partial_{m}^{k}u~dx:=M_{1}.$ (4.16)
First, we expand the nonlinear term using the Leibniz rule and divide it into three parts.
$M_{1}$ $=-2\sum_{m,i,j=1}^{2}\sum_{\alpha=1}^{k}\binom{k}{\alpha}\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}(\partial_{m}^{\alpha}u_{j}\partial_{j}\partial_{m}^{k-\alpha}u_{i})~\Lambda_{1}^{-\sigma}\partial_{m}^{k}u_{i}\,dx$
$=M_{11}+M_{12}+M_{13},$
where the terms $M_{11}$, $M_{12}$, and $M_{13}$ correspond to different choices of the indices $(m,j)$ and are defined as follows. The term $M_{11}$ corresponds to the case $m=1$:
$M_{11}:=\sum_{i,j=1}^{2}\sum_{\alpha=1}^{k}\binom{k}{\alpha}\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}\big(\partial_{1}^{\alpha}u_{j}\,\partial_{j}\partial_{1}^{k-\alpha}u_{i}\big)\,\Lambda_{1}^{-\sigma}\partial_{1}^{k}u_{i}\,dx.$
The term $M_{12}$ corresponds to the case $m=2$ and $j=1$:
$M_{12}:=\sum_{i=1}^{2}\sum_{\alpha=1}^{k}\binom{k}{\alpha}\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}\big(\partial_{2}^{\alpha}u_{1}\,\partial_{1}\partial_{2}^{k-\alpha}u_{i}\big)\,\Lambda_{1}^{-\sigma}\partial_{2}^{k}u_{i}\,dx.$
Finally, the term $M_{13}$ corresponds to the case $m=2$ and $j=2$:
$M_{13}$ $:=\sum_{i=1}^{2}\sum_{\alpha=1}^{k}\binom{k}{\alpha}\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}\big(\partial_{2}^{\alpha}u_{2}\,\partial_{2}\partial_{2}^{k-\alpha}u_{i}\big)\,\Lambda_{1}^{-\sigma}\partial_{2}^{k}u_{i}\,dx$
$=-\sum_{i=1}^{2}\sum_{\alpha=1}^{k}\binom{k}{\alpha}\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}\big(\partial_{1}\partial_{2}^{\alpha-1}u_{1}\,\partial_{2}\partial_{2}^{k-\alpha}u_{i}\big)\,\Lambda_{1}^{-\sigma}\partial_{2}^{k}u_{i}\,dx,$
where we have used the incompressibility condition $\partial_{2}u_{2}=-\partial_{1}u_{1}$.
To estimate each component of $M_{1}$, it suffices to establish the following lemma.
@@
Let $1\leq\beta\leq k$ be an integer. For $\ell\in\{1,2\}$ and $\tfrac{3}{4}\leq s<\sigma+\tfrac{1}{2}$, we have
$\sum_{\beta=1}^{k}\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}\big(\partial^{\beta}_{\ell}u\partial_{1}\partial^{k-\beta}_{\ell}u\big)\Lambda_{1}^{-\sigma}\partial^{k}_{\ell}udx\leq C\|\Lambda_{1}^{-\sigma+s}u\|_{H^{k}}^{2}\|\Lambda_{1}^{-\sigma}u\|_{H^{k}}.$
We now proceed to estimate the term $M_{1}$. The proof of Lemma 4.1 will be given later. Assuming Lemma 4.1 for the moment, we obtain
$|M_{1}|\leq|M_{11}|+|M_{12}|+|M_{13}|\leq C\|\Lambda_{1}^{-\sigma+s}u\|_{H^{k}}^{2}\|\Lambda_{1}^{-\sigma}u\|_{H^{k}}.$
Substituting this bound into (4.2) and integrating in time yield
$\sum_{m=1}^{2}\|\Lambda_{1}^{-\sigma}\partial_{m}^{k}u(t)\|_{L^{2}}^{2}+2\nu\int_{0}^{t}\sum_{m=1}^{2}\|\Lambda_{1}^{s}\Lambda_{1}^{-\sigma}\partial_{m}^{k}u(\tau)\|_{L^{2}}^{2}\,d\tau$
$\leq{}$ $\sum_{m=1}^{2}\|\Lambda_{1}^{-\sigma}\partial_{m}^{k}u_{0}\|_{L^{2}}^{2}+C\int_{0}^{t}\|\Lambda_{1}^{-\sigma+s}u(\tau)\|_{H^{k}}^{2}\|\Lambda_{1}^{-\sigma}u(\tau)\|_{H^{k}}\,d\tau.$ (4.17)
Combining (4.17) with the $L^{2}$ energy estimate (4.14), we arrive at
$\|\Lambda_{1}^{-\sigma}u(t)\|_{H^{k}}^{2}$ $+2\nu\int_{0}^{t}\|\Lambda_{1}^{s}\Lambda_{1}^{-\sigma}u(\tau)\|_{H^{k}}^{2}\,d\tau$
$\leq{}$ $\|\Lambda_{1}^{-\sigma}u_{0}\|_{H^{k}}^{2}+C\int_{0}^{t}\|\Lambda_{1}^{-\sigma+s}u(\tau)\|_{H^{k}}^{2}\|\Lambda_{1}^{-\sigma}u(\tau)\|_{H^{k}}\,d\tau.$ (4.18)

### v2

@@ Abstract
The stability problem for the 2D Navier-Stokes equations with dissipation in only one direction on $\mathbb{R}^{2}$ is not fully understood. This dissipation is in the intermediate regime between the fully dissipative Navier-Stokes and the invisicd Euler. Navier-Stokes solutions in $\mathbb{R}^{2}$ decay algebraically in time while Euler solutions can grow rather rapidly in time. This paper solves the fundamental stability and large-time behavior problem on the anisotropic Navier-Stokes with fractional dissipation $\Lambda_{1}^{2s}$ for all $0\leq s<1$. The case $s=1$ corresponds to the standard one directional dissipation $\partial_{1}^{2}$. Different techniques are developed to treat different ranges of fractional exponents: $0\leq s\leq\frac{3}{4}$, $\frac{3}{4}<s<\frac{11}{12}$, and $\frac{11}{12}\leq s<1$. The final range is the most difficult case, for which we introduce the spatial polynomial $A_{2}$ weights and exploit the boundedness of Riesz transforms on weighted $L^{2}$-spaces.
@@ 1 Introduction
$u_{0}=(u_{1,0},u_{2,0})=\nabla^{\perp}\psi_{0}:=(-\partial_{2}\psi_{0},\partial_{1}\psi_{0}),$
@@
$\|u_{0}\|_{H^{k}}+\|\Lambda_{1}^{-\sigma}\psi_{0}\|_{L^{2}}+\|\Lambda_{1}^{-\sigma}u_{1,0}\|_{L^{2}}+\|\Lambda_{1}^{-\sigma}\partial_{2}u_{1,0}\|_{L^{2}}\leq\varepsilon.$ (1.4)
@@
$\|u(t)\|_{H^{k}}+\nu^{\frac{1}{2}}\left(\int_{0}^{t}\|\Lambda_{1}^{s}u(\tau)\|_{H^{k}}^{2}\,d\tau\right)^{\frac{1}{2}}\leq C\varepsilon,\quad\|\Lambda_{1}^{-\sigma}u\|_{L^{2}}\leq C\varepsilon,\qquad\forall\,t\geq 0,$
@@
A special feature of this result is that the second velocity component $u_{2}$ decays faster than the first component $u_{1}$. This anisotropic decay is revealed through the use of the stream function and is a key ingredient in the proof of the theorem. The enhanced decay of $u_{2}$ plays a crucial role in closing the nonlinear estimates. Moreover, since the Navier-Stokes equations do not propagate negative Sobolev regularity of the form $\Lambda_{1}^{-\sigma}u$ for $\sigma\geq\tfrac{1}{2}$, the restriction $\sigma<\tfrac{1}{2}$ is necessary.
@@
Once again, these estimates shows that the second velocity component $u_{2}$ exhibits a faster decay rate than $u_{1}$, reflecting the underlying anisotropic structure of the system. The derivation of the above estimates parallels that of Theorem 1.2, we only present the estimates for $u_{2}$, which differs slightly, in Appendix B.
@@
$\|u(t)\|_{H^{k}}+\nu^{1/2}\left(\int_{0}^{t}\|\Lambda_{1}^{s}u(\tau)\|_{H^{k}}^{2}\,d\tau\right)^{1/2}\leq C\varepsilon,\quad\|\Lambda_{1}^{-\sigma}u\|_{L^{2}}\leq C\epsilon,\qquad\forall\,t\geq 0.$
@@ 2 Anisotropic Estimates and Tools
A weight function $w\colon\mathbb{R}^{d}\to[0,\infty)$ is said to belong to the $A_{p}$ class if:
@@
    $w\in L_{\mathrm{loc}}^{1}(\mathbb{R}^{d})$ and $w(x)>0$ almost everywhere;
@@
    $\left(\frac{1}{|Q|}\int_{Q}w(y)\,dy\right)\left(\frac{1}{|Q|}\int_{Q}w(y)^{-\frac{1}{p-1}}\,dy\right)^{p-1}\leq C,\qquad 1<p<\infty,$
@@
$N_{1}:=-2\sum_{i,j=1}^{2}\sum_{\alpha=1}^{k}\binom{k}{\alpha}\int_{\mathbb{R}^{2}}\partial_{1}^{\alpha}u_{j}\,\partial_{j}\partial_{1}^{k-\alpha}u_{i}\,\partial_{1}^{k}u_{i}\,dx,$ (3.5)
@@
$N_{2}:=-2\sum_{i=1}^{2}\sum_{\alpha=1}^{k}\binom{k}{\alpha}\int_{\mathbb{R}^{2}}\partial_{2}^{\alpha}u_{1}\,\partial_{1}\partial_{2}^{k-\alpha}u_{i}\,\partial_{2}^{k}u_{i}\,dx,$ (3.6)
@@
$=$  $\;2\sum_{i=1}^{2}\sum_{\alpha=1}^{k}\binom{k}{\alpha}\int_{\mathbb{R}^{2}}\partial_{1}\partial_{2}^{\alpha-1}u_{1}\,\partial_{2}\partial_{2}^{k-\alpha}u_{i}\,\partial_{2}^{k}u_{i}\,dx,$ (3.7)
@@
We decompose the analysis of $F^{III}_{H}$ into two parts according to the cases $1\leq\beta\leq k-1$ and $\beta=k$.
@@
$+\int_{\mathbb{R}^{2}}\varphi(\xi_{1})\mathcal{F}_{x_{1}}\big(\partial^{k}_{\ell}f\partial^{k}_{\ell}f\big)(\xi_{1},x_{2})\mathcal{F}_{x_{1}}\big(\partial_{1}f\big)(\xi_{1},x_{2})d\xi_{1}dx_{2}$
@@
In this section, we prove Theorem 1.2. The argument couples two ingredients: the $H^{k}$ stability estimate for $u$ and the large-time decay estimates for $u$ and its derivatives.
More precisely, by imposing a mild negative Sobolev regularity assumption on the initial data in the horizontal direction, we are able to derive algebraic decay rates for $u$ and $\partial_{1}u$ in $L^{2}$. At the same time, we also prove the propagation of the negative horizontal regularity $\|\Lambda_{1}^{-\sigma}u\|_{L^{2}}$. These decay rates, in turn, render the time integrals of the nonlinear terms arising in the high-order energy estimate convergent, which is the key difficulty when $s>\frac{3}{4}$. Indeed, for $s\leq\frac{3}{4}$ the stability of (1.1) can be established by energy estimates alone (Theorem 1.1), whereas for $s>\frac{3}{4}$ the dissipation $\Lambda_{1}^{2s}$ is no longer strong enough to directly absorb certain nonlinear terms in the $H^{k}$ estimate. The decay of $\|u_{1}\|_{L^{2}}$, $\|\partial_{1}u_{1}\|_{L^{2}}$, $\|\partial_{2}u_{1}\|_{L^{2}}$, $\|u_{2}\|_{L^{2}}$, and $\|\partial_{1}u_{2}\|_{L^{2}}$ is established via Duhamel’s principle, where a key structural observation is that $u_{2}$ decays strictly faster than $u_{1}$ owing to the incompressibility constraint and the stream-function representation. The two estimates are then closed simultaneously through a standard bootstrap argument.
@@
$\|u(t)\|_{H^{k}}$ $\leq\tfrac{1}{2}C_{0}\varepsilon,$ (4.7)
@@
4.1 Estimate of $\|u\|_{H^{k}}$
In this subsection, under the assumption that $u$ and $\partial_{1}u$ enjoy the prescribed time decay, we prove that the solution $u$ satisfies a closed high-order energy estimate in the range $\frac{3}{4}<s<\frac{1}{2}+\sigma$. The main difficulty is the treatment of the nonlinear interactions. By exploiting the decay of $u$ and $\partial_{1}u$, we show that these nonlinear terms can be absorbed into the energy inequality. As a consequence, under the additional assumption that the initial data are sufficiently small, we complete the bootstrap argument for $\|u\|_{H^{k}}$.
We first record the core nonlinear estimate used in the proof.
@@
Assume that $\frac{3}{4}<s<\frac{1}{2}+\sigma$, and that $u$ and $\partial_{1}u$ satisfy the decay bounds postulated in the bootstrap assumptions. Then there exist constants $\vartheta>2$ and $\Omega>1$ such that
$\sum_{\beta=1}^{k}\int_{\mathbb{R}^{2}}\partial_{2}^{\beta}u\,\partial_{1}\partial_{2}^{k-\beta}u\,\partial_{2}^{k}u\,dx\leq C(C_{0}\varepsilon)^{\vartheta}(1+t)^{-\Omega}+C\|\partial_{2}^{k}u\|_{L^{2}}\,\|\Lambda_{1}^{s}\partial_{2}^{k}u\|_{L^{2}}^{2}.$
The proof of this lemma will be given at the end of this subsection.
The $L^{2}$ energy estimate for $u$ is standard:
$\|u(t)\|_{L^{2}}^{2}+2\nu\int_{0}^{t}\|\Lambda_{1}^{s}u(\tau)\|_{L^{2}}^{2}\,d\tau\leq\|u_{0}\|_{L^{2}}^{2}.$ (4.13)
For the $\dot{H}^{k}$ norm, using the same notations as (3.3)–(3.1)), we have
$\frac{d}{dt}\sum_{m=1}^{2}\|\partial_{m}^{k}u\|_{L^{2}}^{2}+2\nu\sum_{m=1}^{2}\|\Lambda_{1}^{s}\partial_{m}^{k}u\|_{L^{2}}^{2}=N=N_{1}+N_{2}+N_{3},$ (4.14)
For $N_{1}$, since there are enough $\partial_{1}$ derivatives, the estimate is direct. Assume $1\leq\alpha\leq k-1$. By Hölder’s inequality and Lemma 2.3, we obtain
$|N_{1}|$ $\leq C\sum_{i,j=1}^{2}\sum_{\alpha=1}^{k-1}\binom{k}{\alpha}\|\partial_{1}^{\alpha}u_{j}\|_{L^{2}_{x_{1}}L^{\infty}_{x_{2}}}\|\partial_{j}\partial_{1}^{k-\alpha}u_{i}\|_{L^{4}_{x_{1}}L^{2}_{x_{2}}}\|\partial_{1}^{k}u_{i}\|_{L^{4}_{x_{1}}L^{2}_{x_{2}}}.$
By Sobolev embedding in the $x_{1}$ variable,
$\|\partial_{j}\partial_{1}^{k-\alpha}u_{i}\|_{L^{4}_{x_{1}}L^{2}_{x_{2}}}\|\partial_{1}^{k}u_{i}\|_{L^{4}_{x_{1}}L^{2}_{x_{2}}}$ $\leq\|\Lambda_{1}^{\frac{1}{4}}\partial_{j}\partial_{1}^{k-\alpha}u_{i}\|_{L^{2}}\|\Lambda_{1}^{\frac{1}{4}}\partial_{1}^{k}u_{i}\|_{L^{2}}$
$\leq\|\Lambda_{1}^{s}\Lambda_{1}^{\frac{1}{4}-s}\partial_{j}\partial_{1}^{k-\alpha}u_{i}\|_{L^{2}}\|\Lambda_{1}^{s}\Lambda_{1}^{\frac{1}{4}-s}\partial_{1}^{k}u_{i}\|_{L^{2}}.$
Substituting this into the above inequality yields
$|N_{1}|$ $\leq C\sum_{i,j=1}^{2}\sum_{\alpha=1}^{k-1}\binom{k}{\alpha}\|\partial_{1}^{\alpha}u_{j}\|_{L^{2}_{x_{1}}L^{\infty}_{x_{2}}}\|\Lambda_{1}^{s}\Lambda_{1}^{\frac{1}{4}-s}\partial_{j}\partial_{1}^{k-\alpha}u_{i}\|_{L^{2}}\|\Lambda_{1}^{s}\Lambda_{1}^{\frac{1}{4}-s}\partial_{1}^{k}u_{i}\|_{L^{2}}$
$\leq C\|u\|_{H^{k}}\|\Lambda_{1}^{s}u\|_{H^{k}}^{2}.$
For the endpoint case $\alpha=k$, we similarly have
$\sum_{i,j=1}^{2}\left|\int_{\mathbb{R}^{2}}\big(\partial_{1}^{k}u_{j}\,\partial_{j}u_{i}\big)\,\partial_{1}^{k}u_{i}\,dx\right|\leq C\|u\|_{H^{k}}\|\Lambda_{1}^{s}u\|_{H^{k}}^{2},$
so the $\alpha=k$ contribution is controlled in the same way.
The new difficulty, compared with Theorem 1.1, lies in $N_{2}$ and $N_{3}$: these terms involve $\partial_{1}$ acting on only one factor, so they cannot be controlled by the dissipation $\Lambda_{1}^{2s}$ alone when $s>\frac{3}{4}$. We handle them using the bootstrap decay bounds (4.2) and (4.5).
Applying Lemma 4.1 to $N_{2}$, and using the identity $\partial_{2}u_{2}=-\partial_{1}u_{1}$, we obtain the same bound for $N_{3}$. Consequently,
$|N_{2}|+|N_{3}|\leq C(C_{0}\varepsilon)^{\vartheta}(1+t)^{-\Omega}+C\|\partial_{2}^{k}u\|_{L^{2}}\,\|\Lambda_{1}^{s}\partial_{2}^{k}u\|_{L^{2}}^{2},$
where $\vartheta>2$ and $\Omega>1$ are given by Lemma 4.1. Since $\|\partial_{2}^{k}u\|_{L^{2}}\leq\|u\|_{H^{k}}$ and $\|\Lambda_{1}^{s}\partial_{2}^{k}u\|_{L^{2}}\leq\|\Lambda_{1}^{s}u\|_{H^{k}}$, we infer that
$|N|\leq C(C_{0}\varepsilon)^{\vartheta}(1+t)^{-\Omega}+C\|u\|_{H^{k}}\|\Lambda_{1}^{s}u\|_{H^{k}}^{2}.$ (4.15)
Substituting (4.15) into (4.14) and integrating with respect to time, we get
$\sum_{m=1}^{2}\|\partial_{m}^{k}u(t)\|_{L^{2}}^{2}+2\nu\int_{0}^{t}\sum_{m=1}^{2}\|\Lambda_{1}^{s}\partial_{m}^{k}u(\tau)\|_{L^{2}}^{2}\,d\tau$
$\leq{}$ $\sum_{m=1}^{2}\|\partial_{m}^{k}u_{0}\|_{L^{2}}^{2}+C(C_{0}\varepsilon)^{\vartheta}+C\int_{0}^{t}\|\Lambda_{1}^{s}u(\tau)\|_{H^{k}}^{2}\|u(\tau)\|_{H^{k}}\,d\tau.$ (4.16)
Combining (4.13) and (4.1), we conclude
$\|u(t)\|_{H^{k}}^{2}+\nu\int_{0}^{t}\|\Lambda_{1}^{s}u(\tau)\|_{H^{k}}^{2}\,d\tau\leq C\|u_{0}\|_{H^{k}}^{2}+C(C_{0}\varepsilon)^{\vartheta}+C\int_{0}^{t}\|\Lambda_{1}^{s}u(\tau)\|_{H^{k}}^{2}\|u(\tau)\|_{H^{k}}\,d\tau.$ (4.17)